OpenStudy (anonymous):

What is the solution set for |–2x + 8| = 12

OpenStudy (amorfide):

there are many ways to do this, what method do you want? graph method? squaring method?

OpenStudy (anonymous):

10 and -2

OpenStudy (goformit100):

I WANT THE CONCEPT

OpenStudy (anonymous):

CaseI: -2x + 8 =12 -2x = 4 x = -2 CaseII: 2x - 8 = 12 2x = 20 x=10

OpenStudy (anonymous):

thanks^

OpenStudy (anonymous):

Welcome.)